Write the equation of a line parallel to the given line but passing through the given point.
y = 1/4x – 4
(8, -1)

Answer by TimothyLamb(4379)

parallel lines have equal slopes
---
slope of new line = (1/4)
---
new line in point-slope form:
y + 1 = (1/4)(x - 8)
---
new line in slope-intercept form:
y + 1 = (1/4)x - 8(1/4)
y = (1/4)x - 2 - 1
y = (1/4)x - 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php